OpenStudy (anonymous):

Antonio is seven less than four times his daughter's age. Antonio is 45; how old is his daughter? Answer 17 15 13 19

OpenStudy (anonymous):

(45+7)/4=13 13(4)=52 52-7=45

OpenStudy (anonymous):

let x be the age of her daughter let 4x-7 be the age of Antonio equation: 4x-7=45 4x=45+7 4x=52 4x/4=52/4 x = 13 checking: 4x-7 = 45 4(13) - 7 = 45 52-7 = 45 45=45